In a group of people 18 blondes, 17 brunettes and 5 redheads. What is the probability of pulling a redhead.
Answer by jim_thompson5910(35256)

There are a total of 18+17+5 = 40 people

5 of them are redheads

-------------------------------------------------------

P(redhead) = (# of redheads)/(# of people total)

P(redhead) = 5/40

P(redhead) = 1/8 (don't forget to reduce fully)


The probability of selecting a redhead is 1/8


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim